If the sum of all the positive divisors of an integer n is equal to 2n, then n is called a perfect number. Which of the following is a perfect number?
(A) 4 (B) 8 (C) 12 (D) 24 (E) 28

Answer by J2R2R(94)

(A) 4 : Factors are 4, 2 and 1. Sum = 7
(B) 8 : Factors are 8, 4, 2 and 1. Sum = 15
(C) 12 : Factors are 12, 6, 4, 3, 2 and 1. Sum = 28
(D) 24 : Factors are 24, 12, 8, 6, 4, 3, 2 and 1. Sum = 60
(E) 28 : Factors are 28, 14, 7, 4, 2 and 1. Sum = 56
Since (E) is the only one where the sum equals twice the original number, 28 is the only perfect number in this list.